OpenStudy (anonymous):

last question for equation... equation sets use digits: 1 to 8

OpenStudy (anonymous):

or 1 to 10

OpenStudy (anonymous):

6*2-3+8=5+4+1+7 6*3+8=(5+4)*2+1+7 6*2+7+1=5+4+3+8 4+6+7+1=2+3+5+8 4+5+2+7=1+3+6+8 1*3*5=8+7-2-4+6

OpenStudy (anonymous):

for 1 to 10 or 1 to 8

OpenStudy (anonymous):

all for 1to8

OpenStudy (anonymous):

k thanks i will ask other question about rabbit run

OpenStudy (anonymous):

have u tried to find the solutions yourself before u ask here?

OpenStudy (anonymous):

yes,but i cant find only few soe...

OpenStudy (anonymous):

i search it about a week

OpenStudy (anonymous):

ok.the trick is u pick a few digits(about half of them or more) that can form an equation.then see where u can put the rest of the digits left.

OpenStudy (anonymous):

okey...i keep it

OpenStudy (anonymous):

or u can put aside a few digits 1st. like my first 4examples, i put aside 7,1,8 bcoz 7+1=8. just manipulate the other digits

OpenStudy (anonymous):

owh....

OpenStudy (anonymous):

can you give conclusions?? :))